# Long-range Ising model for regional-scale seismic risk analysis

### v1

@@ Appendix B Derivation of the mean-field approximation
$x_{i}x_{j}$ $=\left(m_{i}+\left(x_{i}-m_{i}\right)\right)\left(m_{j}+\left(x_{i}-m_{i}\right)\right)$ (19)
@@
Subsequently, the normalizing constant $Z$ can be expressed by:
@@
where $\tilde{h}_{i}=h_{i}+\sum_{j\neq i}J_{ij}m_{j}$ is the effective external field that incorporates the effect of pairwise interactions, and $c\equiv\exp\left(-\frac{1}{2}\sum_{i\neq j}J_{ij}m_{i}m_{j}\right)$ is introduced for brevity. The symmetry of $\boldsymbol{J}$, $J_{ij}=J_{ji}$, was used for the third line of the equation. We further simplify $Z$ into:
@@
$=c\prod_{i=1}^{n}\left(\exp(\tilde{h}_{i})+\exp(-\tilde{h}_{i})\right)$
@@
Finally, using the properties of $m_{i}=\frac{\partial\ln{Z}}{\partial h_{i}}$ and $c_{ij}=\frac{\partial\ln{Z}}{\partial J_{ij}}=\frac{\partial m_{i}}{\partial h_{j}}$, the self-consistent equations for $m_{i}$ and $c_{ij}$ are derived as
@@
$=\tanh\left({h_{i}+\sum_{j\neq i}J_{ij}m_{j}}\right)$

### v2

@@ Appendix B Derivation of the mean-field approximation
The mean-field approximation is foundational in statistical mechanics, and its derivation can be found in many papers and textbooks [33, 34, 60, 18]. However, for completeness, we present a derivation in the context of the long-range Ising model.
@@
$x_{i}x_{j}$ $=\left(m_{i}+\left(x_{i}-m_{i}\right)\right)\left(m_{j}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\left(x_{j}-m_{j}\right)}}\right)$ (19)
@@
Subsequently, the normalizing constant $Z$ can be obtained by summing the unnormalized probabilities in Eq. (9) over all configurations $I$, expressed by:
@@
where $\tilde{h}_{i}=h_{i}+\sum_{j\neq i}J_{ij}m_{j}$ is the effective external field that incorporates the effect of pairwise interactions, and $c\equiv\exp\left(-\frac{1}{2}\sum_{i\neq j}J_{ij}m_{i}m_{j}\right)$ is introduced for brevity. The symmetry of $\boldsymbol{J}$, $J_{ij}=J_{ji}$, was used for the third line of Eq. (20). We further simplify $Z$ into:
@@
$={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}c\sum_{I}\left(\exp{\tilde{h}_{1}x_{1}}\right)\left(\exp{\tilde{h}_{2}x_{2}}\right)\cdots\left(\exp{\tilde{h}_{n}x_{n}}\right)}$
$={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}c\left(\exp\left(\tilde{h}_{1}\right)+\exp\left(-\tilde{h}_{1}\right)\right)\left(\exp\left(\tilde{h}_{2}\right)+\exp\left(-\tilde{h}_{2}\right)\right)\ldots\left(\exp\left(\tilde{h}_{n}\right)+\exp\left(-\tilde{h}_{n}\right)\right)}$
$=c\prod_{i=1}^{n}\left(\exp\left(\tilde{h}_{i}\right)+\exp\left(-\tilde{h}_{i}\right)\right)$
@@
Finally, using the properties of $m_{i}=\frac{\partial\ln{Z}}{\partial h_{i}}$ and $c_{ij}=\frac{\partial m_{i}}{\partial h_{j}}$, the self-consistent equations for $m_{i}$ and $c_{ij}$ are derived as:
@@
$=\tanh\left({h_{i}+\sum_{j\neq i}J_{ij}m_{j}}\right)\,,$